I need help figuring the steps to this problem: The length of a rectangle is five feet less than nine times the width. Find the length, width, and perimeter of the rectangle if the area is 4 square feet.
Answer by rimjhim.tomar(21)

let the width be x feet,
given that length is five feet less than nine times the width.
hence length= 9x-5
now the area of the rectangle is given=4 square feet
length*width=4
x(9x-5)=4
9x^2-5x-4=0
9x^2-9x+4x-4=0
9x(x-1)+4(x-1)=0
(9x+4)(x-1)=0
we get either x=(-4/9) or x=1
but the width of the rectangle can't be negative,
hence x=(-4/5) is invalid.
so we get,
width = x = 1 feet
and length = 9(1)-5 = 4 feet
and perimeter = 2(length + width) = 2(4+1) = 2(5)= 10 feet